John spins a sling around in a circle, preparing to release a projectile with a mass of 28 grams. If his sling is 60 cm long, and he is spinning it at a steady speed - so that the edge of the sling is moving at 30 m/s, what is the centripetal acceleration experienced by the projectile? What is the centripetal force acting on the projectile?

What is the best way to angularly accelerate a door? What is the worst way? Why?

Please answer both of these questions.
Answer #1

Centripetal acceleration = (v×v)/r = 900/0.6 = 1500 m/s/s

Centripetal force= m×a = 0.028× 1500= 42 N

The best way to angularly accelerate a door is to apply the force at the farthest perpendicular distance from the hinge.

Because angular acceleration depends on radius.

The worst way to apply force near the hinge.